OpenStudy (anonymous):

Select the equation for the circle with center at point (0, 5) and a radius of ½.

OpenStudy (lukebluefive):

Equation of a circle: \[(x - h)^2 + (y - k)^2 = radius^2\] Entering given information: \[x^2 + (y - 5)^2 = \left( \frac{1}{2} \right)^2 = \frac{1}{4}\]